to help kim practice for special olympics, aally runs beside him for half the required distance. she runs the remaining distance at her regular speed and arrives 90 secs ahead of kim. what is the ratio of sallys regular speed to kims speed? T= kims total time.


options - 


a) T/90;


b)T/(T-90);


c)T/(T-180);


d)T/180;

Grade:12

1 Answers

Anvitha Reddy
12 years ago

Spped of Kim  K

Speed of Sally S

90* K=T/2*S

   S/K=90*2/T=180/T